# First detection of VHE gamma-ray signal from the FSRQ TON 0599

### v1

@@ 1 Introduction
Blazars are a subclass of active galactic nuclei (AGN) which eject material through collimated jets of ultrarelativistic matter (Urry & Padvani, 1995). They are known for significant and fast flux variability, with time scales as short as order of minutes. Blazars are subdivided into BL Lacertae objects (BL Lacs) and flat spectrum radio quasars (FSRQs). Although BL Lacs generally have featureless optical spectra, FSRQs usually exhibit prominent emission lines and strong continuum emission in the optical-UV band (the so-called blue bump) associated with the emission from the accretion disc. Although FSRQs generally have higher luminosities than BL Lacs, their inverse Compton (IC) bumps peak at lower energies, resulting in softer very high energy (VHE, $E>100$ GeV) spectra. In addition, they are mostly located at higher redshifts (Ackermann et al., 2011), with their VHE spectra further softened by absorption on extragalactic background light (EBL). For these reasons, FSRQs are more difficult to detect in the VHE band, TON 0599 being one of the handful detected so far.
@@
TON 0599 (other names: 4C +29.45, B2 1156+29, TeV J1159+292, RGB J1159+292, RX J1159.5+2914) is located at Right Ascension (J2000.0) 11h59m31.8s and Declination +29d14m44s. At the redshift of $z=0.7247$ (Hewett & Wild, 2010) it is currently the sixth farthest source of VHE gamma rays, and fifth farthest AGN. TON 0599 is a highly variable source in all frequencies (Wills et al., 1983; Wills et al., 1992; Hovatta et al., 2007; Savolainen & Kovalev, 2008; Liu et al., 2013; Wang et al., 2014). Stockman (1978) and Fan et al. (2006) classified it as an optically violently variable (OVV) quasar based on polarization and photometric measurements, respectively. In addition, Fan et al. (2006) found possible periodicity in the optical light curve with time scales of $1.58$ and $3.55$ years. The flux variability in near-infrared was followed in 2007–2013 at the Campo Imperatore and Teide observatories, monitored by the GLAST-AGILE Support Program (GASP) of the WEBT. The fractional variability (see Section 3.2) in J band in that period was 0.73 (Raiteri et al., 2014). Also, intranight variability at the level of 22.3 per cent of the flux was observed with Sampurnanand Telescope, possibly caused by turbulence in the jet (Goyal et al., 2013).
@@
The first detection of TON 0599 in gamma rays was reported in the second EGRET catalogue (Thompson et al., 1995). It was also listed in the First Fermi Large Area Telescope (LAT) source catalogue (Abdo et al., 2010). TON 0599 underwent a particulary active flaring state in late 2017. Multiwavelength (MWL) studies analysing that period were published in Patel et al. (2018); Prince (2019); Hallum et al. (2022); Patel & Chitnis (2019).
@@
On 2017 December 15, TON 0599 was detected for the first time in the VHE band with the Major Atmospheric Gamma Imaging Atmospheric (MAGIC) telescopes (Mirzoyan, 2017). The details of the detection are reported in this work. In Section 2, we describe the observations used in this study and lay out the results. The MWL picture is presented in Section 3, and the results of the broadband modelling in Section 3.3. We summarise our results in Section 4.
@@ 2.1.1 MAGIC observation and analysis
MAGIC is a system of two 1717\,m mirror dish Imaging Atmospheric Cherenkov Telescopes (IACTs) located in El Roque de los Muchachos observatory in Canary Island of La Palma ($28.7^{\circ}$ N, $17.9^{\circ}$ W), at the altitude of 2200 m a.s.l. (Aleksić et al., 2016a). MAGIC observations of TON 0599 started on 2017 December 15 (MJD 58102), following a hardening of the HE gamma-ray spectrum observed with Fermi-LAT. The observations lasted until 2017 December 29 (MJD 58116). During this time, 19.2 hours of data were collected. All observations were performed in the so-called wobble mode (Fomin et al., 1994; Aleksić et al., 2016b), under dark sky conditions (Aleksić et al., 2016a). Based on measurements with a pyrometer and an AllSky camera (Will, 2017) and subjective estimation by the observers [^1] The MAGIC LIDAR system used for monitoring the atmospheric transmission (Fruck & Gaug, 2015) was not functioning properly during the campaign., the atmospheric transmission was estimated to be above 85 per cent with respect to clean atmosphere for all observations. The data were reduced and analysed using MAGIC Analysis and Reconstruction Software (MARS, Zanin et al. 2013). For points whose relative uncertainties are greater than 0.5, the upper limits on the flux have been computed following Rolke et al. (2005) at the confidence level of 95 per cent.
On the first night, after observing for 0.96 hours with the MAGIC telescopes, the signal was detected with a statistical significance of 10.4 $\sigma$ (using Eq. 17 of Li & Ma 1983). The integral flux above 80 GeV reached $(2.6\pm 0.3)\times 10^{-10}$ cm-2 s-1, corresponding to 42 per cent of flux from the Crab Nebula (0.42 C.U.) in the same energy range, as reported in Aleksić et al. (2015). The second night showed a marginal increase in flux (although within uncertainties) to $(3.1\pm 0.3)\times 10^{-10}$ cm-2 s-1, which corresponds to 0.51 C.U. above 80 GeV. After the second night, the measured flux decreased until the end of the MAGIC observation campaign. The light curve above 80 GeV is shown in Fig. 1, while in Fig. 6 it is put in the MWL context.
@@ 2.1.3 Gamma-ray results
where $E_{0}$ is the normalisation energy fixed here to 95 GeV, $f_{0}$ the normalisation flux, $\alpha$ the spectral slope, and $E_{c}$ the cut-off energy. The fitted SED is represented by the dashed red line in Fig. 2. The fit results are given in Table 2, where we report on statistical uncertainties only. The systematic uncertainty on the MAGIC energy scale is 15 per cent (Aleksić et al., 2016c). The systematic uncertainty on the spectral slope of the MAGIC spectra in all four periods is estimated to be $\pm 0.15$; and the systematic uncertainty on the cut-off energy is $15$ per cent. For the Fermi-LAT data, the uncertainties were dominated by statistical ones, which were taken into account when fitting the spectra. Given that MAGIC systematics dominate over Fermi-LAT ones, we estimated the resulting systematic uncertainty of the spectral slope of the joint fit at $\pm 0.10$, making the total systematic uncertainty on the slope $\pm 0.18$.
@@ 2.2.2 Swift/XRT
The Swift/XRT operates in the 0.3100.3-10 keV energy range. The XRT spectra were obtained from the processed event files available through the UK Swift Science Data Centre online data products generator (Evans et al., 2009). For each observing period, time-averaged spectra were extracted and analysed using the HEASOFT v6.20 software package. The spectral files were grouped using the grppha task to ensure a minimum of 20 counts per bin, allowing the use of $\chi^{2}$ statistics. Spectral fitting was performed with XSPEC v12.9.1 (Dorman & Arnaud, 2001). The spectra were modeled with an absorbed power-law model. Photoelectric absorption was taken into account using a fixed hydrogen column density corresponding to the Galactic value along the line of sight, $N_{\text{H}}=1.77\times 10^{20}$ cm-2 (Willingale et al., 2013). For each period, the absorbed power-law model provides a good description of the data. From the spectral fits, we derived the photon index $\alpha$ and the unabsorbed flux in the 0.3100.3-10 keV energy band. The best-fit spectral parameters are reported in Table 4. The flux appears to be rather high, but constant at the beginning of the extended observation period. Between MJD 58056 and 58076, the average flux value is about 20 per cent higher than the highest flux observed during the MAGIC flare. This coincides partially with the highest flux in Fermi-LAT, although because of the sparse X-ray sampling and the lack of observations in other wavebands, it is impossible to determine the origin of the high X-ray flux and whether it follows the variability visible in the HE gamma-ray band. Later, during the MAGIC flare, the highest X-ray flux is lower by some 25 per cent, compared to the highest flux overall, and it decays towards the end of the MAGIC flare. The X-ray spectrum plays a fundamental role in constraining the emission models, since the X-ray region corresponds to the valley between the synchrotron and the high-energy bump. As can be seen in Fig. 8, the XRT flux contributes to the high-energy peak.
@@ 2.3 WEBT
Radio observations were collected as part of the WEBT campaign. They were performed at the Noto and Medicina (Italy, Giroletti & Righini, 2020), Metsähovi (Finland, Teraesranta et al., 1998), Pico Veleta (Spain, IRAM 30 m telescope) observatories and at the Submillimeter Array (SMA, Maunakea, USA). Measurements from the IRAM 30 m Telescope were obtained as part of the POLAMI (Polarimetric Monitoring of AGN at Millimetre Wavelengths) Program [^7] http://polami.iaa.es (Agudo et al., 2018b; Agudo et al., 2018a; Thum et al., 2018). The SMA data are from the SMA Calibrator Flux Density database [^8] http://sma1.sma.hawaii.edu/callist/callist.html (Gurwell et al., 2007). The radio light curves are shown in Fig. 6.
@@ 3.2 Fractional variability
Schleicher et al. (2019) discussed how the completeness of datasets affects fractional variability and found that the relative systematic uncertainties due to sampling can be as large as 10 percent. This could particularly influence the comparison between the variabilities in the two time intervals. However, in our case, the higher variability for the longer period is most likely due to large differences in fluxes over a longer time interval. Moreover, as Schleicher et al. (2019) themselves argue, the systematic uncertainties are too small to affect the conclusions on the general trends.
@@ 3.3 SED modelling
We model the emission from TON 0599 using the simple one-zone leptonic model fully described in Maraschi & Tavecchio (2003). In this scheme, the low-energy peak is modeled as synchrotron emission from relativistic electrons carried by the jet. The high-energy peak is attributed to IC from the same electrons (Ghisellini et al., 1998). In FSRQs the dominant target photon population is generally assumed to be external, i.e. the radiation from the broad line region (BLR) and/or the dusty torus (e.g. Ghisellini & Tavecchio 2009). We consider a homogeneous blob of relativistic electrons moving down the jet with bulk Lorentz factor $\Gamma_{b}$ and carrying a tangled magnetic field with (comoving) strength $B^{\prime}$. In the frame of the emission region the electrons follow a smoothed broken power-law energy distribution extending from $\gamma^{\prime}_{\mathrm{min}}$ to $\gamma^{\prime}_{\mathrm{max}}$. Spectral indices below and above the break at $\gamma^{\prime}_{\mathrm{brk}}$ are $p_{1}$ and $p_{2}$, respectively, and the normalisation is $K^{\prime}$ (all primed quantities are expressed in the source frame). We model emission in each period separately and independently. We consider the emission in each period as static, i.e. we do not model transition from one period to the next.
@@
According to Ghisellini & Tavecchio (2009) the radii of the BLR and dust torus are proportional to the square root of total emission of the accretion disk. The disk luminosity is estimated from the BLR emission lines luminosity based on a recipe from Ghisellini & Tavecchio (2015) to be $3.6\times 10^{45}$ erg s-1, which is similar, although below the lower limit used in Patel et al. (2018). This gives $1.9\times 10^{17}$ cm and $4.7\times 10^{18}$ cm for the radii of the BLR and dust torus, respectively. Given the prominent broad line emission, a strong absorption of gamma rays would be expected if the gamma-ray emission zone was within the BLR zone (see e.g. Donea & Protheroe 2003; Sitarek & Bednarek 2008; Tavecchio & Mazin 2009; Poutanen & Stern 2010; Costamante et al. 2018). Indeed we do detect a cut-off in the spectra between Fermi-LAT and MAGIC bands (see Table 2), however at energies significantly higher than the expected $\sim 20$ GeV, and the $\sim 12$ GeV reported by Patel & Chitnis (2019). Therefore, as suggested for the case of other FSRQ showing the same behaviour (see e.g. Aleksić et al. 2011, 2014; Ahnen et al. 2015, 2016), we place the emission region outside of the BLR, setting a lower limit on the emission zone distance from the black hole (BH) $d>1.9\times 10^{17}$ cm. This estimate of the BLR radius is consistent with the values previously reported in Patel et al. (2018, $(2.11-2.45)\times 10^{17}$ cm) or Prince (2019, $(2.4-2.98)\times 10^{17}$ cm). Assuming the jet aperture angle $\theta_{j}=0.1$ rad, one obtains a lower limit on the radius of the emission zone of $R>1.9\times 10^{16}$ cm. We set the upper limits on the gamma-ray emission zone based on the causality argument
@@
where $c$ is the speed of light, $\delta$ is the relativistic Doppler factor, and $z$ redshift of the source. For the variability timescale $t_{\mathrm{var}}$, we take the flux halving time from Table 6. Using $t_{\mathrm{var}}=3.9$ days and $\delta=20$ (intermediate value from our modelling reported in Table 8), we obtain $R\leq 1.2\times 10^{17}$ cm. The final emission zone radius is obtained from the model fit, and the distance from the BH is fixed at 10 times the value. By construction, the only important external radiation field is the one produced by the torus (see e.g. Ghisellini & Tavecchio 2009), modelled as a black body at 1000 K.
The results of the modelling are shown in Fig. 8, and the inferred model parameters are reported in Table 8. The parameters are found by judging “by eye” the agreement between the model output and the observational data points. Moreover, we tried to minimise the variations of the parameters between the different states, and the two slopes of the electron energy distribution have been fixed. Since the models corresponding to states A and B are very similar, in Fig. 8 and Table 8 we only report the model and the parameters corresponding to state A. The values that we got for some of the parameters are substantially different from the ones reported in Patel et al. (2018) (e.g. we got a factor of $\sim 10-30$ larger emission zone radius, while a factor of $\sim 30-80$ weaker magnetic field); however, Patel et al. (2018) base their estimates on data in X-ray and lower energy bands and fit the SED with a second degree polynomial, while our results are based on broadband modelling across the whole electromagnetic spectrum and using a physical scenario. In our model synchrotron radiation accounts for the low-energy bump, up to the ‘valley’ between the two bumps. synchrotron self-Compton (SSC) contributes substantially (about 50 per cent) only in the X-ray band (see Fig. 8, where we separately report the contributions of the different emission components for state D). Above the X-ray band, the emission is dominated by external Compton (EC) scattering on the torus photons. The different states are reproduced with slight variations of some of the model parameters, while the main difference between states is a result of the variation of the bulk Lorentz factor, which decreases with time. However, in view of the degeneracies characterising the emission model, it is possible that other combinations of the parameters can satisfactorily reproduce the data.
@@
Considering $\Gamma_{b}\geq 10$ (as we get from our emission modelling consistently with standard modelling, e.g. Ghisellini & Tavecchio 2009), one would expect that the blob would travel for large distances on a time-scale of 10 days. However, a presence of a stationary (in observer’s frame) shock, in which the plasma flows at relativistic speed (see e.g. Bodo & Tavecchio 2018) could explain high bulk Lorentz factors without the emission region covering large distances. The same scenario would explain lower flux levels corresponding to lower values of $\Gamma_{b}$.
This scenario does not exclude some degree of dynamical behaviour. For instance, Bodo & Tavecchio (2018) studied the effect of a density perturbation on the structure of the recollimation shock and found that the structure of the shock (that, in turn, influences the value of physical quantities in the downstream region, where the emission occurs) shows important variations. In particular, the change of the shock shape (and therefore of the inclination of the upstream velocity with respect to the shock front) determines strong variations of the downstream velocity pattern (both in direction and intensity) that can strongly affect the inferred effective Lorentz factor and size used in our simplified one-zone model. A detailed study of the feasibility of the proposed scenario is beyond the scope of the present article.
@@
Electron normalisation [cm-3] $K^{\prime}$ $3.0\times 10^{3}$ $3.9\times 10^{3}$ $3.7\times 10^{3}$

### v2

@@ 1 Introduction
Blazars are a subclass of active galactic nuclei (AGN) which eject material through collimated jets of ultrarelativistic matter (Urry and Padovani, 1995). They are known for significant and fast flux variability, with time scales as short as order of minutes. Blazars are subdivided into BL Lacertae objects (BL Lacs) and flat spectrum radio quasars (FSRQs). Although BL Lacs generally have featureless optical spectra, FSRQs usually exhibit prominent emission lines and strong continuum emission in the optical-UV band (the so-called blue bump) associated with the emission from the accretion disc. Although FSRQs generally have higher luminosities than BL Lacs, their inverse Compton (IC) bumps peak at lower energies, resulting in softer very high energy (VHE, $E>100$ GeV) spectra. In addition, they are mostly located at higher redshifts (Ackermann et al., 2011), with their VHE spectra further softened by absorption on extragalactic background light (EBL). For these reasons, FSRQs are more difficult to detect in the VHE band, TON 0599 being one of the handful detected so far.
@@
TON 0599 (other names: 4C +29.45, B2 1156+29, TeV J1159+292, RGB J1159+292, RX J1159.5+2914) is located at Right Ascension (J2000.0) 11h59m31.8s and Declination +29d14m44s. At the redshift of $z=0.7247$ (Hewett and Wild, 2010) it is currently the sixth farthest source of VHE gamma rays, and fifth farthest AGN. TON 0599 is a highly variable source in all frequencies (Wills et al., 1983, 1992; Hovatta et al., 2007; Savolainen and Kovalev, 2008; Liu et al., 2013; Wang et al., 2014). Stockman (1978) and Fan et al. (2006) classified it as an optically violently variable (OVV) quasar based on polarization and photometric measurements, respectively. In addition, Fan et al. (2006) found possible periodicity in the optical light curve with time scales of $1.58$ and $3.55$ years. The flux variability in near-infrared was followed in 2007–2013 at the Campo Imperatore and Teide observatories, monitored by the GLAST-AGILE Support Program (GASP) of the Whole Earth Blazar Telescope (WEBT) Collaboration. The fractional variability (see Section 3.2) in J band in that period was 0.73 (Raiteri et al., 2014). Also, intranight variability at the level of 22.3 per cent of the flux was observed with Sampurnanand Telescope, possibly caused by turbulence in the jet (Goyal et al., 2013).
@@
The first detection of TON 0599 in gamma rays was reported in the second EGRET catalogue (Thompson et al., 1995). It was also listed in the First Fermi Large Area Telescope (LAT) source catalogue (Abdo et al., 2010). TON 0599 underwent a particulary active flaring state in late 2017. Multiwavelength (MWL) studies analysing that period were published in Patel et al. (2018); Prince (2019); Hallum et al. (2022); Patel and Chitnis (2019).
@@
On 2017 December 15, TON 0599 was detected for the first time in the VHE band with the Major Atmospheric Gamma Imaging Atmospheric (MAGIC) telescopes (Mirzoyan, 2017). The details of the detection are reported in this work. In Section 2, we describe the observations used in this study and lay out the results. The MWL picture is presented in Section 3, and the results of the broadband modelling in Section 3.3. We summarize our results in Section 4.
@@ 2.1.1 MAGIC observation and analysis
MAGIC is a system of two 1717\,m mirror dish Imaging Atmospheric Cherenkov Telescopes (IACTs) located in El Roque de los Muchachos observatory in Canary Island of La Palma ($28.7^{\circ}$ N, $17.9^{\circ}$ W), at the altitude of 2200 m a.s.l. (Aleksić et al., 2016b). MAGIC observations of TON 0599 started on 2017 December 15 (MJD 58102), following a hardening of the HE gamma-ray spectrum observed with Fermi-LAT. The observations lasted until 2017 December 29 (MJD 58116). During this time, 19.2 hours of data were collected. All observations were performed in the so-called wobble mode (Fomin et al., 1994; Aleksić et al., 2016a), under dark sky conditions (Aleksić et al., 2016b). Based on measurements with a pyrometer and an AllSky camera (Will, 2017) and subjective estimation by the observers [^1] The MAGIC LIDAR system used for monitoring the atmospheric transmission (Fruck and Gaug, 2015) was not functioning properly during the campaign., the atmospheric transmission was estimated to be above 85 per cent with respect to clean atmosphere for all observations. The data were reduced and analysed using MAGIC Analysis and Reconstruction Software (MARS, Zanin et al. 2013). For points whose relative uncertainties are greater than 0.5, the upper limits on the flux have been computed following Rolke et al. (2005) at the confidence level of 95 per cent.
On the first night, after observing for 0.96 hours with the MAGIC telescopes, the signal was detected with a statistical significance of 10.4 $\sigma$ (using Eq. 17 of Li and Ma 1983). The integral flux above 80 GeV reached $(2.6\pm 0.3)\times 10^{-10}$ cm-2 s-1, corresponding to 42 per cent of flux from the Crab Nebula (0.42 C.U.) in the same energy range, as reported in Aleksić et al. (2015). The second night showed a marginal increase in flux (although within uncertainties) to $(3.1\pm 0.3)\times 10^{-10}$ cm-2 s-1, which corresponds to 0.51 C.U. above 80 GeV. After the second night, the measured flux decreased until the end of the MAGIC observation campaign. The light curve above 80 GeV is shown in Fig. 1, while in Fig. 6 it is put in the MWL context.
@@ 2.1.3 Gamma-ray results
where $E_{0}$ is the normalization energy fixed here to 95 GeV, $f_{0}$ the normalization flux, $\alpha$ the spectral slope, and $E_{c}$ the cut-off energy. The fitted SED is represented by the dashed red line in Fig. 2. The fit results are given in Table 2, where we report on statistical uncertainties only. The systematic uncertainty on the MAGIC energy scale is 15 per cent (Aleksić and others, 2016). The systematic uncertainty on the spectral slope of the MAGIC spectra in all four periods is estimated to be $\pm 0.15$; and the systematic uncertainty on the cut-off energy is $15$ per cent. For the Fermi-LAT data, the uncertainties were dominated by statistical ones, which were taken into account when fitting the spectra. Given that MAGIC systematics dominate over Fermi-LAT ones, we estimated the resulting systematic uncertainty of the spectral slope of the joint fit at $\pm 0.10$, making the total systematic uncertainty on the slope $\pm 0.18$.
@@ 2.2.2 Swift/XRT
The Swift/XRT operates in the 0.3100.3-10 keV energy range. The XRT spectra were obtained from the processed event files available through the UK Swift Science Data Centre online data products generator (Evans et al., 2009). For each observing period, time-averaged spectra were extracted and analysed using the HEASOFT v6.20 software package. The spectral files were grouped using the grppha task to ensure a minimum of 20 counts per bin, allowing the use of $\chi^{2}$ statistics. Spectral fitting was performed with XSPEC v12.9.1 (Dorman and Arnaud, 2001). The spectra were modeled with an absorbed power-law model. Photoelectric absorption was taken into account using a fixed hydrogen column density corresponding to the Galactic value along the line of sight, $N_{\text{H}}=1.77\times 10^{20}$ cm-2 (Willingale et al., 2013). For each period, the absorbed power-law model provides a good description of the data. From the spectral fits, we derived the photon index $\alpha$ and the unabsorbed flux in the 0.3100.3-10 keV energy band. The best-fit spectral parameters are reported in Table 4. The flux appears to be rather high, but constant at the beginning of the extended observation period. Between MJD 58056 and 58076, the average flux value is about 20 per cent higher than the highest flux observed during the MAGIC flare. This coincides partially with the highest flux in Fermi-LAT, although because of the sparse X-ray sampling and the lack of observations in other wavebands, it is impossible to determine the origin of the high X-ray flux and whether it follows the variability visible in the HE gamma-ray band. Later, during the MAGIC flare, the highest X-ray flux is lower by some 25 per cent, compared to the highest flux overall, and it decays towards the end of the MAGIC flare. The X-ray spectrum plays a fundamental role in constraining the emission models, since the X-ray region corresponds to the valley between the synchrotron and the high-energy bump. As can be seen in Fig. 8, the XRT flux contributes to the high-energy peak.
@@ 2.3 WEBT
Radio observations were collected as part of the WEBT campaign. They were performed at the Noto and Medicina (Italy, Giroletti and Righini, 2020), Metsähovi (Finland, Teraesranta et al., 1998), Pico Veleta (Spain, IRAM 30 m telescope) observatories and at the Submillimeter Array (SMA, Maunakea, USA). Measurements from the IRAM 30 m Telescope were obtained as part of the POLAMI (Polarimetric Monitoring of AGN at Millimetre Wavelengths) Program [^7] http://polami.iaa.es (Agudo et al., 2018a, b; Thum et al., 2018). The SMA data are from the SMA Calibrator Flux Density database [^8] http://sma1.sma.hawaii.edu/callist/callist.html (Gurwell et al., 2007). The radio light curves are shown in Fig. 6.
@@ 3.2 Fractional variability
Schleicher and others (2019) discussed how the completeness of datasets affects fractional variability and found that the relative systematic uncertainties due to sampling can be as large as 10 percent. This could particularly influence the comparison between the variabilities in the two time intervals. However, in our case, the higher variability for the longer period is most likely due to large differences in fluxes over a longer time interval. Moreover, as Schleicher and others (2019) themselves argue, the systematic uncertainties are too small to affect the conclusions on the general trends.
@@ 3.3 SED modelling
We model the emission from TON 0599 using the simple one-zone leptonic model fully described in Maraschi and Tavecchio (2003). In this scheme, the low-energy peak is modeled as synchrotron emission from relativistic electrons carried by the jet. The high-energy peak is attributed to IC from the same electrons (Ghisellini et al., 1998). In FSRQs the dominant target photon population is generally assumed to be external, i.e. the radiation from the broad line region (BLR) and/or the dusty torus (e.g. Ghisellini and Tavecchio 2009). We consider a homogeneous blob of relativistic electrons moving down the jet with bulk Lorentz factor $\Gamma_{b}$ and carrying a tangled magnetic field with (comoving) strength $B^{\prime}$. In the frame of the emission region the electrons follow a smoothed broken power-law energy distribution extending from $\gamma^{\prime}_{\mathrm{min}}$ to $\gamma^{\prime}_{\mathrm{max}}$. Spectral indices below and above the break at $\gamma^{\prime}_{\mathrm{brk}}$ are $p_{1}$ and $p_{2}$, respectively, and the normalization is $K^{\prime}$ (all primed quantities are expressed in the source frame). We model emission in each period separately and independently. We consider the emission in each period as static, i.e. we do not model transition from one period to the next.
@@
According to Ghisellini and Tavecchio (2009) the radii of the BLR and dust torus are proportional to the square root of total emission of the accretion disk. The disk luminosity is estimated from the BLR emission lines luminosity based on a recipe from Ghisellini and Tavecchio (2015) to be $3.6\times 10^{45}$ erg s-1, which is similar, although below the lower limit used in Patel et al. (2018). This gives $1.9\times 10^{17}$ cm and $4.7\times 10^{18}$ cm for the radii of the BLR and dust torus, respectively. Given the prominent broad line emission, a strong absorption of gamma rays would be expected if the gamma-ray emission zone was within the BLR zone (see e.g. Donea and Protheroe 2003; Sitarek and Bednarek 2008; Tavecchio and Mazin 2009; Poutanen and Stern 2010; Costamante et al. 2018). Indeed we do detect a cut-off in the spectra between Fermi-LAT and MAGIC bands (see Table 2), however at energies significantly higher than the expected $\sim 20$ GeV, and the $\sim 12$ GeV reported by Patel and Chitnis (2019). Therefore, as suggested for the case of other FSRQ showing the same behaviour (see e.g. Aleksić et al. 2011, 2014; Ahnen et al. 2015, 2016), we place the emission region outside of the BLR, setting a lower limit on the emission zone distance from the black hole (BH) $d>1.9\times 10^{17}$ cm. This estimate of the BLR radius is consistent with the values previously reported in Patel et al. (2018, $(2.11-2.45)\times 10^{17}$ cm) or Prince (2019, $(2.4-2.98)\times 10^{17}$ cm). Assuming the jet aperture angle $\theta_{j}=0.1$ rad, one obtains a lower limit on the radius of the emission zone of $R>1.9\times 10^{16}$ cm. We set the upper limits on the gamma-ray emission zone based on the causality argument
@@
where $c$ is the speed of light, $\delta$ is the relativistic Doppler factor, and $z$ redshift of the source. For the variability timescale $t_{\mathrm{var}}$, we take the flux halving time from Table 6. Using $t_{\mathrm{var}}=3.9$ days and $\delta=20$ (intermediate value from our modelling reported in Table 8), we obtain $R\leq 1.2\times 10^{17}$ cm. The final emission zone radius is obtained from the model fit, and the distance from the BH is fixed at 10 times the value. By construction, the only important external radiation field is the one produced by the torus (see e.g. Ghisellini and Tavecchio 2009), modelled as a black body at 1000 K.
The results of the modelling are shown in Fig. 8, and the inferred model parameters are reported in Table 8. The parameters are found by judging “by eye” the agreement between the model output and the observational data points. Moreover, we tried to minimize the variations of the parameters between the different states, and the two slopes of the electron energy distribution have been fixed. Since the models corresponding to states A and B are very similar, in Fig. 8 and Table 8 we only report the model and the parameters corresponding to state A. The values that we got for some of the parameters are substantially different from the ones reported in Patel et al. (2018) (e.g. we got a factor of $\sim 10-30$ larger emission zone radius, while a factor of $\sim 30-80$ weaker magnetic field); however, Patel et al. (2018) base their estimates on data in X-ray and lower energy bands and fit the SED with a second degree polynomial, while our results are based on broadband modelling across the whole electromagnetic spectrum and using a physical scenario. In our model synchrotron radiation accounts for the low-energy bump, up to the ‘valley’ between the two bumps. synchrotron self-Compton (SSC) contributes substantially (about 50 per cent) only in the X-ray band (see Fig. 8, where we separately report the contributions of the different emission components for state D). Above the X-ray band, the emission is dominated by external Compton (EC) scattering on the torus photons. The different states are reproduced with slight variations of some of the model parameters, while the main difference between states is a result of the variation of the bulk Lorentz factor, which decreases with time. However, in view of the degeneracies characterising the emission model, it is possible that other combinations of the parameters can satisfactorily reproduce the data.
@@
Considering $\Gamma_{b}\geq 10$ (as we get from our emission modelling consistently with standard modelling, e.g. Ghisellini and Tavecchio 2009), one would expect that the blob would travel for large distances on a time-scale of 10 days. However, a presence of a stationary (in observer’s frame) shock, in which the plasma flows at relativistic speed (see e.g. Bodo and Tavecchio 2018) could explain high bulk Lorentz factors without the emission region covering large distances. The same scenario would explain lower flux levels corresponding to lower values of $\Gamma_{b}$.
This scenario does not exclude some degree of dynamical behaviour. For instance, Bodo and Tavecchio (2018) studied the effect of a density perturbation on the structure of the recollimation shock and found that the structure of the shock (that, in turn, influences the value of physical quantities in the downstream region, where the emission occurs) shows important variations. In particular, the change of the shock shape (and therefore of the inclination of the upstream velocity with respect to the shock front) determines strong variations of the downstream velocity pattern (both in direction and intensity) that can strongly affect the inferred effective Lorentz factor and size used in our simplified one-zone model. A detailed study of the feasibility of the proposed scenario is beyond the scope of the present article.
@@
Electron normalization [cm-3] $K^{\prime}$ $3.0\times 10^{3}$ $3.9\times 10^{3}$ $3.7\times 10^{3}$